12123333333333333333:

A marine biologist is studying the growth of a particular species of fish. She writes the following equation to show the length of the fish, f(m), in cm, after m months: f(m) = 3(1.09)m Part A: When the marine biologist concluded her study, the length of the fish was approximately 5.98 cm. What is a reasonable domain to plot the growth function? (4 points) I believe the answer is 0 < x < 9 Part B: What does the y-intercept of the graph of the function f(m) represent? (2 points) The y-intercept (3) represents how big the fish was when she first started her study (at 0 months). Part C: What is the average rate of change of the function f(m) from m = 2 to m = 8, and what does it represent? (4 points) idk

12123333333333333333:

@Gdeinward

12123333333333333333:

I believe I have everything correct but idk for part c..im working on it.

Gdeinward:

so does f(m) = 3(1.09)m have any exponents?

12123333333333333333:

f(m)=3(1.09)^m

12123333333333333333:

I think the average rate of change is 0.309.....I think

12123333333333333333:

Is that correct?

Gdeinward:

well to find ARC, you divide y1 - y2 / x1 - x2

Gdeinward:

so ehen m equals 2, m is the x variable, and your resulting number after you solve the equation is your y variable

Gdeinward:

so solve f(m) = 3(1.09)^2 to find ur y then solve f(m) = 3(1.09)^8 to find your second y

Gdeinward:

DO you understand?

12123333333333333333:

So 0.309?

12123333333333333333:

:P I think it is 0.309! i DO undertand

12123333333333333333:

oof.

Gdeinward:

eh I hadnt done the math, give me a sec

12123333333333333333:

Ok

Gdeinward:

I got 0.6475145

12123333333333333333:

hm

12123333333333333333:

How did u do it?

12123333333333333333:

I used the formula in my school

12123333333333333333:

\[\frac{ 3-4.235}{ 0-4}\]

Gdeinward:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Gdeinward so solve f(m) = 3(1.09)^2 to find ur y then solve f(m) = 3(1.09)^8 to find your second y \(\color{#0cbb34}{\text{End of Quote}}\) I solved the above equations to get f(m) = 3(1.09)^2 = f(m) = 3.885087 f(m) = 3(1.09)^8 = f(m) = 5.97768792507

Gdeinward:

then I subtracted answer a from answer B to get 2.09260792507

12123333333333333333:

f(b)−f(a) over b−a